[Hallc_running] [New Logentry] Hall C RC Update

latif at jlab.org latif at jlab.org
Thu Mar 14 10:20:04 EDT 2019

Logentry Text:
Current Status
- Yesterday kinematic group-36 was completed around 7:00 PM.

- During the transition to kinematic group-37, we changed momentum setting for both spectrometers including polarity change for SHMS. The spectrometer rotation required an escorted-access to reset the drive controller (logbook entry:https://logbooks.jlab.org/entry/3667412).

- During the owl shift, the loop 3 heater power went to zero without any alarm (logbook entry: https://logbooks.jlab.org/entry/3667554). It was restored after turning off the beam and adjusting the JT valve.

- We are currently taking data for kinematic setting 37-8. 

Future Plan 

- We will perform a BCM calibration around 10:00 AM today.

- Once done with kinematic group 37, we will move to the last kinematic group 38.

- CSV will be over around 4:00 PM on Sunday. In preparation for installation on Monday, we will ramp down all the HMS and SHMS magnets, rotate the HMS to about 40 degrees, rotate the SHMS to about 20 degrees. target expert will come in to warm up the target in the afternoon.

- Detector and beam pipe installation work will take place starting Sunday swing through Tuesday day shift.

- We informed MCC that we are expecting 3rd pass setup at the end of day shift on Tuesday.

- Shifts starting from Sunday (March 17th) swing shift through Tuesday (March 18th) day shifts are canceled. Shift takers on Tuesday Swing shift are on standby and should check e-mail prior to coming to the counting-house. 

- There is no update to the plan discussed during yesterday's RC meeting. So we will NOT have any RC meeting today.


- Today's run plan: https://hallcweb.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Wednesday_March_14
- Online analysis instruction: https://hallcweb.jlab.org/wiki/index.php/CSV_Fall_2018_Run_Plan
- Information for Shift Takers: https://hallcweb.jlab.org/wiki/index.php/Information_for_Shift_Takers


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3667709
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190314/4ccef53e/attachment.html>

More information about the Hallc_running mailing list